ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 1 Eyles, Joseph W. ,University of Texas at Austin, USA Discovering R. L. Moore’s Calculus Course. R. L. Moore (1882-1974) was distinguished professor of mathematics at the University of Texas from 1920 until 1969. While the Moore Method is described by noted mathematician and author Paul Halmos as the “right way to teach anything and everything, ” his pedagogical techniques are described in the context of graduate level and upper division mathematics courses. Although the fact that Dr. Moore taught calculus every year is often mentioned in the literature, it is not clear how he employed his techniques in this typically non-theorem-proving course. A description of the calculus course taught by Dr. Moore will be presented. Reference: Paul R. Halmos, I Want To Be A Mathematician, The Mathematical Association of America, Washington, D. C., 1985.